myininaya (myininaya):

i think the problem was sin(2x+3)=cos(30-x) and we want to find the minimal value of x using the factor theorem does this look right joe?

myininaya (myininaya):

i found a solution but i don't how to use the factor theorem like that one guy asked for

OpenStudy (anonymous):

yeah thats right, is there something more to this problem? i thought we solved it?

myininaya (myininaya):

he wanted minimal value for x

myininaya (myininaya):

and i have no clue how to use factor theorem to solve this

OpenStudy (anonymous):

oh oh. i think by factor theorem he meant sum/difference formulas, it would be a long drawn out process >.<

myininaya (myininaya):

is that what you were attempting?

OpenStudy (anonymous):

yes, and i got tired lol >.<

myininaya (myininaya):

lol so i got x=pi/2-75

OpenStudy (anonymous):

your solution was swift and fast, like a ninja.

myininaya (myininaya):

but i don't know how i'm suppose to show its minimal value

myininaya (myininaya):

there are other solutions

OpenStudy (anonymous):

i guess i could do it my way and see what i get >.< give me an hour lol

OpenStudy (anonymous):

Is the angles in degrees here

myininaya (myininaya):

its in radians

OpenStudy (anonymous):

inside the sin was 2x+45 tw

OpenStudy (anonymous):

im sry, 2x-45

OpenStudy (anonymous):

i remember negative root 2 over 2's popping up

myininaya (myininaya):

roshan this is what i did earlier \[\sin(2x+45)=\sin(\frac{\pi}{2}-(30-x))\] \[2x+45=\frac{\pi}{2}-(30-x)\]

myininaya (myininaya):

thanks joe

myininaya (myininaya):

\[2x-x=\frac{\pi}{2}-30-45\] \[x=\frac{\pi}{2}-75\]

myininaya (myininaya):

trig functions are funky how do you know you have all the solutions

myininaya (myininaya):

and i know i don't

OpenStudy (anonymous):

this is too long, i dont wanna do it >.<

myininaya (myininaya):

ok don't do it lol

OpenStudy (anonymous):

your solution looks funny o.O are those numbers in degrees?

myininaya (myininaya):

no

myininaya (myininaya):

lol thats what rosahan asked lol

myininaya (myininaya):

we also have the solution \[x=-\frac{3\pi}{2}-75\]

OpenStudy (anonymous):

75 almost like 12(2pi)....its like your flying around the unit circle redundantly o.O

OpenStudy (anonymous):

if that is indeed radians.

myininaya (myininaya):

i also checked it out wolfram to see if the solution worked and it did i think we can come up with something not so ugly

myininaya (myininaya):

maybe not i'm not sure

myininaya (myininaya):

\[x=\frac{5\pi}{2}-75 \] is another solution i want to generalize my answer

myininaya (myininaya):

oh i see how to :)

OpenStudy (anonymous):

whats the pattern?

myininaya (myininaya):

\[x=(-1)^{n}(2n+1)\frac{\pi}{2}-75, n=0,1,2,3,4,5,6...\]

myininaya (myininaya):

now since this goes on and on there is no way to tell what is the smallest solution

OpenStudy (anonymous):

Somehow my teacher told us to change the whole thing to sin theta and move it to one side.